What mass of calcium chloride will be produced from 3.750 moles of HCl reacting with Ca(OH)2?
1
Expert's answer
2020-05-03T15:22:43-0400

Ca(OH)2 + 2HCl = CaCl2 + 2H2O

According to the equation

n(CaCl2) = n(HCl)/2 = 3.75/2 = 1.875 mol

m(CaCl2) = n*M = 1.875mol*111g/mol = 208.125 g